# Personalized Collaborative Fine-Tuning for On-Device Large Language Models

### v1

@@ 1 Introduction
  • We demonstrate the inefficacy of uniform averaging in personalized LLMs and present 3 unique collaboration protocols for identifying optimal collaborators. Our results showcase superior performance compared to local fine-tuning.
@@ 2 Related Works
We focus on a fully decentralized learning setting, where the existence of a central server is not assumed. Instead, the end users conduct peer-to-peer communication using decentralized schemes such as gossip averaging to aggregate local information across agents. To enable a personalized weighted aggregation, several recent works have proposed data-dependent communication protocols as with task similarities and node qualities. Notably, Zhang et al. (2021) derives a first-order approximation for optimal aggregation weights $w_{ij}^{\star}$, which happens to be proportional to how well client $j$’s model generalizes on client $i$’s data. Li et al. (2022) directly optimizes the mixing weights by minimizing the local validation loss per node. Sui et al. (2022) uses the E-step of the EM algorithm to gauge the significance of other agents for a specific agent $i$, achieved through assessing the accuracy of these agents’ models on the local data of agent $i$. Fan et al. (2024) adopt a different approach by comparing prediction similarity on a publicly available dataset, which largely reduces communication overhead and allows for model heterogeneity. These approaches have been showing promising performances in supervised deep learning experiments. In the language domain, the data heterogeneity across users is not well studied, and the effectiveness of the collaborator selection methods has not been explored for LLMs, where self-supervised training is performed.
@@ 3 Methods
We work in the setting where all users have the same LoRA rank $r$, and aim to arrive at their personalized $\Delta{\bm{\Phi}}_{i}$. During the fine-tuning phase, only LoRA parameters will be updated, and the original weights of the pre-trained model will remain frozen. Therefore, it suffices to communicate LoRA weights, instead of the big chunk of whole parameters.
@@ 3.1 Our Protocol
We stick to the standard model updates sharing and aggregation approach. Each user, through their unique data distribution, contributes to the collective learning process by sharing local model updates. Central to our approach is the so-called trust matrix, which we define as the gossip aggregation matrix to guide each user’s gradient aggregation. The calculation of trust matrix ${\bm{W}}$ will be detailed in Section 3.2. A simplified diagram of our protocol is shown in Figure 1.
@@
Figure 1: Diagram of our protocol. ${\bm{\theta}}_{i}$ are the LoRA weights, and $\Delta{\bm{\theta}}_{i}$ are LoRA weight updates. $s_{i}$ denotes messages to send beside $\Delta{\bm{\theta}}_{i}$. $g(\cdot)$ is our proposed trust calculation method.
@@ 3.2 Trust Calculation
In this section, we explore different ways of building trust – gossip aggregation graph across users. Essentially, the weight of edge $(i,j)$ should denote to what extent can user $j$ ’s gradients help to facilitate user $j$’s learning progress.
@@ 3.2.1 Weights similarity based
$\tilde{w}_{ij}^{t}=\frac{\langle{\bm{\theta}}_{i}^{(t-1)},{\bm{\theta}}_{j}^{(t-1)}\rangle}{\|{\bm{\theta}}_{i}^{(t-1)}\|\|{\bm{\theta}}_{j}^{(t-1)}\|}$ (2)
${\bm{W}}^{t}=\text{SoftMax}(\widetilde{{\bm{W}}}^{t},\text{dim}=1)$
@@ 3.2.2 Validation performance based
Drawing inspiration from Zhang et al. (2021), we evaluate the pairwise trust $w_{ij}$ based on how well user $i$’s model performs on the validation sets of user $j$. This method requires extra validation sets within each user, on which the performance assessment directly reflects the relevance and potential benefit of one user’s model to another. The aim is to guide the aggregation process by favoring models that demonstrate compatible performance levels, thereby enhancing the overall effectiveness of collaborative learning. The calculation is denoted as:
$\tilde{w}_{ij}^{t}=\mathcal{L}(f_{{\bm{\theta}}_{i}^{t-1}}({\bm{X}}_{j}^{\text{val}}),{\bm{X}}_{j}^{\text{val}})$ (3)
${\bm{W}}^{t}=\text{SoftMax}(-\widetilde{{\bm{W}}}^{t},\text{dim}=1)$
where $f_{{\bm{\theta}}_{i}^{t-1}}({\bm{X}}_{j}^{\text{val}})$ denotes a forward pass of user $i$’s model on user $j$’s local validation set, and $\mathcal{L}$ is the cross entropy loss for next token prediction.
@@ 3.2.3 Prediction similarity based
$\tilde{w}_{ij}^{t}=|f_{{\bm{\theta}}_{i}^{t-1}}({\bm{X}}_{S})-f_{{\bm{\theta}}_{j}^{t-1}}({\bm{X}}_{S})|$ (4)
${\bm{W}}^{t}=\text{SoftMax}(-\widetilde{{\bm{W}}}^{t},\text{dim}=1)$
@@ 3.3 Remarks
Table 1: A comparison of communication and computation complexity across the strategies. $L$ denotes the number of layers, $d$ denotes the embedding size, $n$ denotes context length and $k$ denotes LoRA rank. $N$ denotes the number of users in the system. It is clear strategy 2 adds the most extra computational cost, and strategy 3 the second. Strategy 1 requires the least amount of computation on top.
@@ 4.1.1 Experimental details
Our experimental setup was structured as follows: Each user was configured with uniform model architecture, to ensure model updates can be shared and aggregated. We chose relatively small LLMs constrained by the limited computing resources in academia. For each user, we equip them with a GPT2111https://github.com/karpathy/nanoGPT base model, with 124 million parameters in total. It has 12 layers, 12 attention heads, embedding size 768 and vocabulary size 50257.
Figure 2: Ablation study of whether to add LoRA modules compared to full fine-tuning. Left bars correspond to Ratio, which denotes the decrease in validation perplexity compared to the pre-trained model per trainable parameter (higher is better); right bars correspond to perplexity, which denotes validation performance after fine-tuning (lower is better).
LoRA modules were applied on top of both Causal Self-Attention (SA) and MLP layers, constituting $0.47\%$ of full parameters. The reason why we chose the two blocks is a trade-off between the model performance and communication overhead, as shown in Figure 2. Although incorporating LoRA modules onto Self Attention layers results in the most significant reduction in validation perplexity per trainable parameter, integrating LoRA modules into MLP layers can contribute to further decreasing validation perplexity while maintaining a reasonable number of trainable parameters.
@@ 4.1.2 Datasets
Datasets # users # Categories # Training tokens # Validation tokens
@@
Table 2: Dataset configurations used in the experiments. Reference data ${\bm{X}}_{S}$ is slightly larger in size than validation tokens and $\sim 100^{\prime}000$ tokens are sampled for logits computation. There is not enough data available for AG News to establish a meaningful public reference dataset.
@@ 4.2 Results
Our main findings are summarized in Table 4. Apart from local fine-tuning and FedAvg, we incorporate a strong baseline – theoretical. The determination of theoretical collaboration weights relies on the similarity of underlying data source distributions222for example, if user 1 has 1/4 German texts and 3/4 French texts, and user 2 has 1/4 German texts and 3/4 Italian texts, the theoretical weights would be determined by the dot product of $[1/4,3/4,0]$ and $[1/4,0,3/4]$. After obtaining the pairwise dot product matrix, row normalization is performed., which is typically unknown in practice. When compared to two baseline approaches – local training and FedAvg – our validation performance and prediction-based aggregation methods consistently outperform across all datasets and scenarios. Among them, the predictions-based method exhibits the most superior validation performance. Notably, the weights similarity-based aggregation method yields performance akin to FedAvg, where the trust weight is uniformly set to $1/N$ across all clients.
Our methods can sometimes outperform theoretical aggregation, suggesting that a dynamic collaborator selection protocol might be favored in different fine-tuning stages.
@@
Strategy 3 - 36.70(0.23) 17.35(0.18)
Theoretical $\mathit{29.56(0.21)}$ $\mathit{39.55(0.19)}$ $\mathit{17.42(0.19)}$
@@
Strategy 3 - 36.92(0.16) 16.23(0.12)
Theoretical $\mathit{28.08(0.11)}$ $\mathit{35.96(0.24)}$ $\mathit{16.20(0.05)}$
Table 4: Validation perplexities (standard deviation) of our proposed strategies and baseline methods (lower is better). Strategy 1: weights similarity-based; strategy 2: validation performance-based, strategy 3: prediction similarity-based.
@@ 4.3 Trust matrix
We compare our learned trust matrices with the theoretical trust matrix in Figures 3 and 4. It is evident that strategies 2 and 3 uncover similar collaboration patterns as suggested by the theoretical matrix, whereas strategy 1 falls short in this regard. Specifically, strategy 1 assigns nearly identical trust weights to all other users, indicating nearly identical LoRA weights learned at convergence. This strange behavior is further investigated in Appendix A.3: even without communication, the LoRA weights across users are almost equally different, suggesting that LoRA weights are not informative in collaborator selection when cosine similarity is used as the distance metric. This could be attributed to the high-dimensional, sparse nature, and small magnitudes of LoRA weights. The success of strategies 2 and 3 underscores the advantage of utilizing predictions to identify collaborators. It is interesting how the distinctions between different users’ models become more pronounced after a forward pass. Such a phenomenon was not previously observed in the vision domain.
(a) Theoretical
@@
Figure 3: theoretical trust matrix versus learned trust matrix using strategy 1 when users are allocated with Multilingual Wikipedia datasets. The diagonal entries are masked out and the trust is measured when the training converges.
@@
Figure 4: Learned trust matrix using strategies 2 and 3 when users are allocated with Multilingual Wikipedia datasets. The diagonal entries are masked out and the trust is measured when the training converges.
@@ 5 Conclusions
Drawing from insights in collaborative learning communities, we propose 3 collaboration protocols tailored for jointly on-device LLM fine-tuning. To our best knowledge, this is the first study on personalized collaborative LLM fine-tuning. These methods can efficiently address challenges encountered in personalized LLMs, such as limited availability of local data and data heterogeneity stemming from user characteristics.
@@ A.1 Communication cost experiments
Table 5: A comparison of communication complexity across the different shared units. $L$ denotes the number of layers, $d$ denotes the embedding size, $k$ is the LoRA rank, $V$ denotes the vocabulary size and $K$ denotes number of logits kept for communication. $COO$ denotes the sparse encoding used, Coordinate list.
@@ A.2 Hyper-parameters details
The choice of our hyper-parameters is influenced by resource constraints, primarily due to memory limitations (training conducted on a single NVIDIA A100 GPU with 40GB memory) and time constraints, as we needed to fine-tune up to 9 large language models (LLMs). Through empirical observation, we found that a batch size of 50 with 4 accumulation steps and a context length of 512 performed well within memory constraints. Additionally, while increasing the LoRA rank slightly improved validation perplexity, the associated increase in training time rendered it uninteresting for our experiments. Therefore, we opted for LoRA rank 4. Three hyper-parameters remain to be tuned: learning rate, LoRA alpha, and LoRA dropout. In Figure 5, the best perplexity was achieved with a learning rate of 0.002, LoRA dropout of 0.1, and LoRA alpha of 32.
@@
A.3 Weight similarity based trust further analysis

### v2

@@ 1 Introduction
Following some computational social science literatures (Urena et al., 2019; Zhang et al., 2022), we term the aggregation weight matrix “trust”, as if we view clients as members of a social group, their exchanged information can thus be regarded as opinions. Similarity measures can be useful for assessing trust between users, as users are more likely to consider and trust others with similar opinions. Such similarity measures are core to our proposed protocols, which we will discuss in Section 3.1.
@@
  • We demonstrate that collaboration can be leveraged to improve personalization performance in language modeling and present 3 unique collaboration protocols for identifying optimal collaborators.
@@ 2 Related Works
We focus on a peer-to-peer decentralized learning setting, where the existence of a central server is not assumed. Instead, the end users conduct peer-to-peer communication using decentralized schemes such as gossip averaging to aggregate local information across agents. To enable a personalized weighted aggregation, several recent works have proposed data-dependent communication protocols based on task similarities and node qualities. Notably, Zhang et al. (2021) derives a first-order approximation for optimal aggregation weights $w_{ij}^{\star}$, which happens to be proportional to how well client $j$’s model generalizes on client $i$’s data. Li et al. (2022) directly optimizes the mixing weights by minimizing the local validation loss per node. Sui et al. (2022) uses the E-step of the EM algorithm to gauge the significance of other agents for a specific agent $i$, achieved through assessing the accuracy of these agents’ models on the local data of agent $i$. Fan et al. (2024) adopt a different approach by comparing prediction similarity on a publicly available dataset, which largely reduces communication overhead and allows for model heterogeneity. These approaches have been showing promising performances in supervised deep learning experiments. In the language domain, the data heterogeneity across users is not well studied, and the effectiveness of the collaborator selection methods has not been explored for LLMs, where self-supervised training is performed.
@@ 3 Methods
We work in a setting where all users have the same LoRA rank $r$, and aim to arrive at their personalized $\Delta{\bm{\Phi}}_{i}$. During the fine-tuning phase, only LoRA parameters will be updated, and the original weights of the pre-trained model will remain frozen. Therefore, it suffices to communicate LoRA weights, instead of the big chunk of whole parameters.
@@ 3.1 Our Protocol
We stick to the standard approach of sharing and aggregating model updates. Each user, through their unique data distribution, contributes to the collective learning process by sharing local model updates. Central to our approach is the so-called trust matrix, which we define as the gossip aggregation matrix to guide each user’s gradient aggregation. The calculation of trust matrix ${\bm{W}}$ will be detailed in Section 3.2. A simplified diagram of our protocol is shown in Figure 1.
@@
Figure 1: Diagram of our protocol. ${\bm{\theta}}_{i}$ and $\Delta{\bm{\theta}}_{i}$ represent LoRA weights and LoRA weight updates respectively. $s_{i}$ denotes messages to send beside $\Delta{\bm{\theta}}_{i}$, which represents either ${\bm{\theta}}_{i}$ or $f_{{\bm{\theta}}_{i}}({\bm{X}}_{S})$ depending on the protocol (see Table 1). $g(\cdot)$ is our proposed trust calculation approach as detailed in Section 3.2.
@@ 3.2 Trust Calculation
In this section, we explore different ways of building trust – gossip aggregation graph across users. Essentially, the weight of edge $(i,j)$ should denote to what extent can user $j$’s gradients help to facilitate user $i$’s learning progress.
@@ 3.2.1 Weights similarity based
$\tilde{w}_{ij}^{t}=\frac{\langle{\bm{\theta}}_{i}^{(t-1)},{\bm{\theta}}_{j}^{(t-1)}\rangle}{\|{\bm{\theta}}_{i}^{(t-1)}\|\|{\bm{\theta}}_{j}^{(t-1)}\|},\qquad{\bm{W}}^{t}=\text{SoftMax}(\widetilde{{\bm{W}}}^{t},\text{dim}=1)$ (2)
@@ 3.2.2 Validation performance based
Drawing inspiration from Zhang et al. (2021), we evaluate the pairwise trust $w_{ij}$ based on how well user $j$’s model performs on the validation sets of user $i$. This method requires extra validation sets within each user, on which the performance assessment directly reflects the relevance and potential benefit of one user’s model to another. The aim is to guide the aggregation process by favoring models that demonstrate compatible performance levels, thereby enhancing the overall effectiveness of collaborative learning. The calculation is denoted as:
$\tilde{w}_{ij}^{t}=\mathcal{L}(f_{{\bm{\theta}}_{j}^{t-1}}({\bm{X}}_{i}^{\text{val}}),{\bm{X}}_{i}^{\text{val}}),\qquad{\bm{W}}^{t}=\text{SoftMax}(-\widetilde{{\bm{W}}}^{t},\text{dim}=1)$ (3)
where $f_{{\bm{\theta}}_{j}^{t-1}}({\bm{X}}_{i}^{\text{val}})$ denotes a forward pass of user $j$’s model on user $i$’s local validation set, and $\mathcal{L}$ is the cross entropy loss for next token prediction.
@@ 3.2.3 Prediction similarity based
$\tilde{w}_{ij}^{t}=|f_{{\bm{\theta}}_{i}^{t-1}}({\bm{X}}_{S})-f_{{\bm{\theta}}_{j}^{t-1}}({\bm{X}}_{S})|,\qquad{\bm{W}}^{t}=\text{SoftMax}(-\widetilde{{\bm{W}}}^{t},\text{dim}=1)$ (4)
@@ 3.3 Remarks
Table 1: A comparison of communication and computation complexity across the strategies. $L$ denotes the number of layers, $d$ denotes the embedding size, $n$ denotes context length and $k$ denotes LoRA rank. $N$ denotes the number of users in the system. It is clear strategy 2 adds the most extra computational cost, followed by strategy 3. Strategy 1 requires the least amount of computation on top.
@@ 4.1.1 Experimental details
Our experimental setup was structured as follows: Each user was configured with a uniform model architecture, to ensure model updates can be shared and aggregated. We chose relatively small LLMs constrained by the limited computing resources in academia. For each user, we equip them with a GPT2111https://github.com/karpathy/nanoGPT base model, with 124 million parameters in total. It has 12 layers, 12 attention heads, embedding size 768 and vocabulary size 50257.
Figure 2: Ablation study of whether to add LoRA modules compared to full fine-tuning. Left bars correspond to Ratio, which denotes the decrease in test perplexity compared to the pre-trained model per trainable parameter (higher is better); right bars correspond to perplexity, which denotes test performance after fine-tuning (lower is better).
LoRA modules were applied on top of both Causal Self-Attention (SA) and MLP layers, constituting $0.47\%$ of full parameters. The reason why we chose the two blocks is a trade-off between the model performance and communication overhead, as shown in Figure 2. Although incorporating LoRA modules onto Self Attention layers results in the most significant reduction in test perplexity per trainable parameter, integrating LoRA modules into MLP layers can contribute to further decreasing test perplexity while maintaining a reasonable number of trainable parameters.
@@ 4.1.2 Datasets
Datasets # users # Categories # Training tokens # Test tokens
@@
Table 2: Dataset configurations used in the experiments. Reference data ${\bm{X}}_{S}$ is slightly larger in size than test tokens and $\sim 100^{\prime}000$ tokens are sampled for logits computation. There is insufficient data from AG News to establish a meaningful public reference dataset, so we resort to ChatGPT to generate a synthetic one.
@@
4.1.3 Baselines
Our chosen baseline methods are: 1) Local Fine-Tuning, where there is no communication between users, 2) FedAvg (McMahan et al., 2017), where model updates are aggregated to obtain a global model for each user, and 3) FedAvg + Local Fine-Tuning (Jiang et al., 2023), where we spare 10% of training data for local fine-tuning after FedAvg training. We offer an extended table where we vary the amount of training data used for fine-tuning for completeness in Table 9 in Appendix. We further incorporate a strong baseline – oracle. The determination of oracle collaboration weights relies on the similarity of underlying data source distributions, which is typically unknown in practice. We give an example of how the oracle weights are determined: if user 1 has 1/4 German texts and 3/4 French texts, and user 2 has 1/4 German texts and 3/4 Italian texts, the oracle weights would be determined by the dot product of $[1/4,3/4,0]$ and $[1/4,0,3/4]$. After obtaining the pairwise dot product matrix, row normalization is performed.
@@ 4.2 Results
Our main findings are summarized in Table 5. Compared to baseline approaches—local training, FedAvg, and FedAvg + FT—our test performance and prediction-based aggregation methods consistently outperform them across all datasets and scenarios. We emphasize that in language modeling, achieving a balance between personalization and collaboration is not straightforward. Despite significant differences between categories, there are notable similarities in sentence structure and word choices, alongside domain-specific vocabulary. The collaboration-then-personalization approach (FedAvg + Local Fine-Tuning) fails to achieve this balance. However, our methods successfully accomplish it.
Among our protocols, the predictions-based method exhibits the most superior test performance. Notably, the weights similarity-based aggregation method yields performance akin to FedAvg, where the trust weight is uniformly set to $1/N$ across all clients. Our methods can sometimes outperform oracle aggregation, suggesting that a dynamic collaborator selection protocol might be favored in different fine-tuning stages.
Ablation studies on the effect of LoRA ranks and local training set size can be found in Table 8 and Table 7 in Appendix A.3. Our prediction-based strategy consistently delivers superior performance.
A note on communication topology:
While the paper demonstrates a fully connected communication topology, our protocols can support sparser communication patterns, as long as the graph represented by ${\bm{W}}$ is strongly connected (i.e., every vertex is reachable from every other vertex), though convergence will be slower. We provide experimental results on the training time required for a ring topology to achieve the same perplexity level as a fully connected topology in Table 4. In high heterogeneity scenarios where each client is allocated a specific category, a ring topology results in cases where adjacent clients do not share any categories, leading to an oracle trust weight of 0 and hence the NA entries. Despite this, our algorithm can still learn, converging to a perplexity approximately one point higher than reported for the fully connected case.
Method Multi-Wiki Codes-Wiki
Low High Low High
Local 1 1 1 1
FedAvg 1.6 4.2 2.0 1.7
Strategy 1 5.9 12.0 2.0 2.1
Strategy 2 1.1 2.6 1.5 NA
Strategy 3 2.9 NA 1.2 3.8
Table 4: Training time ($X$ times the needed training iterations as in the fully connected case) required for a ring topology to achieve the same perplexity level as a fully connected topology. NA indicates we did not reach the same perplexity after ten times the training iterations.
Figure 3: L1 norm of differences between ${\bm{W}}$ at two consecutive steps for Strategy 2 and Strategy 3
@@
Strategy 3 29.81(0.13) 36.70(0.23) 17.35(0.18)
Oracle $\mathit{29.56(0.21)}$ $\mathit{39.55(0.19)}$ $\mathit{17.42(0.19)}$
@@
Strategy 3 28.72(0.27) 36.92(0.16) 16.23(0.12)
Oracle $\mathit{28.08(0.11)}$ $\mathit{35.96(0.24)}$ $\mathit{16.20(0.05)}$
Table 5: Test perplexities (standard deviation) of our proposed strategies and baseline methods (lower is better). Strategy 1: weights similarity-based; strategy 2: validation performance-based, strategy 3: prediction similarity-based. $\star$ denotes that the used public dataset is a synthetic dataset we constructed using ChatGPT.
@@ 4.3 Trust matrix
We compare our learned trust matrices with the oracle trust matrix in Figures 4 and 5. It is evident that strategies 2 and 3 uncover similar collaboration patterns as suggested by the oracle matrix, whereas strategy 1 falls short in this regard. Specifically, strategy 1 assigns nearly identical trust weights to all other users, indicating nearly identical LoRA weights learned at convergence. This strange behavior is further investigated in Appendix A.4: even without communication, the LoRA weights across users are almost equally different, suggesting that LoRA weights are not informative in collaborator selection when cosine similarity is used as the distance metric.
The success of strategies 2 and 3 underscores the advantage of utilizing predictions to identify collaborators. It is interesting how the distinctions between different users’ models become more pronounced after a forward pass. Such a phenomenon was not previously observed in the vision domain. We further noticed that trust matrices in strategy 3 are more stable across communication rounds compared to strategy 2, with fewer abrupt changes in values. As shown in Figure 3, the $l_{1}$ norm of the differences between ${\bm{W}}$s at consecutive steps fluctuates significantly for Strategy 2. It is speculated that the more stable trust matrix in strategy 3 contributes to stabilizing the learning process, leading to better results.
(a) Oracle
@@
Figure 4: Oracle trust matrix versus learned trust matrix using strategy 1 when users are allocated with Multilingual Wikipedia datasets. The diagonal entries are masked out and the trust is measured when the training ends.
@@
Figure 5: Learned trust matrix using strategies 2 and 3 when users are allocated with Multilingual Wikipedia datasets. The diagonal entries are masked out and the trust is measured when the training ends.
@@ 5 Conclusions
Drawing from insights in collaborative learning communities, we propose 3 collaboration protocols tailored for jointly on-device LLM fine-tuning. To our best knowledge, this is the first work demonstrating that collaboration can be leveraged to improve personalization performance in language modeling. Our methods can efficiently address challenges encountered in personalized LLMs, such as limited availability of local data and data heterogeneity stemming from user characteristics.
@@ 6 Discussions and Future Work
Acknowledgments
We thank Soumajit Majumder and Bettina Messmer for their helpful review and all the anonymous reviewers for their constructive feedback. We acknowledge funding from Huawei Cloud Intelligent Cloud Technologies Initiative, Google Research Collaborations, and from the Swiss National Science Foundation (SNSF) grant number 200020_200342.
@@ A.1 Communication cost experiments
Table 6: A comparison of communication complexity across the different shared units. $L$ denotes the number of layers, $d$ denotes the embedding size, $k$ is the LoRA rank, $V$ denotes the vocabulary size and $K$ denotes the number of logits kept for communication. $COO$ denotes the sparse encoding used, Coordinate list.
@@ A.2 Hyper-parameters details
The choice of our hyper-parameters is influenced by resource constraints, primarily due to memory limitations (training conducted on a single NVIDIA A100 GPU with 40GB memory) and time constraints, as we needed to fine-tune up to 9 large language models (LLMs). Through empirical observation, we found that a batch size of 50 with 4 accumulation steps and a context length of 512 performed well within memory constraints. Additionally, while increasing the LoRA rank slightly improved test perplexity, the associated increase in training time rendered it uninteresting for our experiments. Therefore, we opted for LoRA rank 4. Three hyper-parameters remain to be tuned: learning rate, LoRA alpha, and LoRA dropout. In Figure 6, the best perplexity was achieved with a learning rate of 0.002, LoRA dropout of 0.1, and LoRA alpha of 32.
@@
A.3 Additional Ablations
A.3.1 Impact of local dataset size
We increase the number of training tokens by 100 times, resulting in 84 million training tokens per client in the multilingual setup. The ranking of all methods remained consistent with what we report in Table 5.
Method Local FedAvg Strategy 1 Strategy 2 Strategy 3 Oracle
val ppl 37.80(0.15) 56.69(0.35) 48.47(2.55) 36.17(0.30) 35.63(0.15) 35.59 (0.16)
Table 7: Comparison of methods on Multi-Wiki dataset
A.3.2 Impact of LoRA ranks
In the main body, we restrict the LoRA rank to 4 to maintain a low computational budget. We conduct an ablation study on the LoRA ranks and present the results in Table 8. The rankings of the methods are consistent with our previous findings, with strategy 3 being the best method. Additionally, as expected, a higher LoRA rank results in lower test perplexity.
LoRA Ranks 4 8 16 32
Local 40.24(0.25) 33.15(0.15) 29.14(0.11) 27.41(0.11)
FedAvg 53.23(0.51) 48.84(0.34) 42.68(0.06) 40.12(0.25)
Strategy 1 49.34(2.46) 48.37(1.26) 40.50(0.11) 34.65(5.73)
Strategy 2 37.20(0.20) 30.41(0.13) 26.36(0.12) 24.37(0.20)
Strategy 3 36.92(0.12) 30.25(0.19) 26.30(0.11) 24.39(0.02)
Oracle 35.96(0.24) 30.06(0.16) 26.16(0.07) 24.43(0.10)
Table 8: Comparison of different methods with various LoRA ranks
A.3.3 FedAvg with Local Fine-Tuning
To understand better the relationship between the performance of FedAvg + Local Fine-Tuning and the relative size of the local fine-tuning dataset, we spare different percentages of local data for fine-tuning purposes, and the results are presented in Table 9. With more local fine-tuning data, the performance gets closer to Local training without collaboration.
Heterogeneity Method Datasets
AG News Multi-Wiki Codes-Wiki
Low Local $30.17(0.17)$ $40.00(0.33)$ $19.57(0.23)$
FedAvg $31.66(0.20)$ $52.75(0.57)$ $17.53(0.19)$
FedAvg +FT (10%) $32.25(0.20)$ $48.02(0.25)$ $20.17(0.33)$
FedAvg +FT (25%) $31.29(0.15)$ $43.74(0.24)$ $20.10(0.25)$
FedAvg +FT (50%) $30.03(0.23)$ $39.79(0.21)$ $19.59(0.26)$
Strategy 1 $31.43(0.35)$ $45.59(0.66)$ $17.57(0.21)$
Strategy 2 29.75(0.23) $36.93(0.17)$ $17.61(0.40)$
Strategy 3 29.81(0.13) 36.70(0.23) 17.35(0.18)
Oracle $\mathit{29.56(0.21)}$ $\mathit{39.55(0.19)}$ $\mathit{17.42(0.19)}$
High Local $28.67(0.13)$ $40.24(0.25)$ $17.56(0.08)$
FedAvg $32.08(0.13)$ $53.23(0.51)$ $16.68(0.06)$
FedAvg +FT (10%) $33.74(0.17)$ $48.07(0.18)$ $18.43(0.18)$
FedAvg +FT (25%) $32.11(0.16)$ $43.68(0.13)$ $18.15(0.11)$
FedAvg +FT (50%) $29.84(0.09)$ $40.14(0.19)$ $17.88(0.16)$
Strategy 1 $31.93(0.86)$ $49.34(2.46)$ $16.84(0.05)$
Strategy 2 28.29(0.06) $37.20(0.20)$ $16.22(0.17)$
Strategy 3 28.72(0.27) 36.92(0.16) 16.23(0.12)
Oracle $\mathit{28.08(0.11)}$ $\mathit{35.96(0.24)}$ $\mathit{16.20(0.05)}$
Table 9: (Extended) Test perplexities (standard deviation) of our proposed strategies and baseline methods (lower is better). Strategy 1: weights similarity-based; strategy 2: validation performance-based, strategy 3: prediction similarity-based. $\star$ denotes that the used public dataset is a synthetic dataset we constructed using ChatGPT.
A.4 Weight similarity based trust further analysis